OpenStudy (anonymous):

!?!? How do you even do this without a calculator? square root "a"/ 1+square root "a"

OpenStudy (anonymous):

Better view of the equation \[\sqrt{a}/1+\sqrt{a}\]

OpenStudy (anonymous):

a^1/2 --------- right ? 1 + a^1/2

OpenStudy (anonymous):

no \[(\sqrt{a})/ (1+\sqrt{a)}\]

hero (hero):

multiply top and bottom by (1 - sqrt (a)) then simplify

OpenStudy (anonymous):

what should i do here ...the main deal in these question is to have no root in denominator so what you can do is multiply and divide by 1 - rada rad a is same as a^1/2

OpenStudy (anonymous):

the answer the book gave is : \[(\sqrt{a}-a)/ (1-\sqrt{a})\]

OpenStudy (anonymous):

I am not sure on how to do that Hero :/

hero (hero):

Yes, exactly...That's what you get when you multiply top and bottom by 1 - sqrt(a) , then simplify

hero (hero):

Okay, I can show you

hero (hero):

I will post a classroom

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

oh? what is that?

hero (hero):

You'll see

OpenStudy (anonymous):

ohh..Sorry I dont have a mic or webcam, will it still work?

hero (hero):

Yes

OpenStudy (anonymous):

sorry my bandwith is poor

myininaya (myininaya):

\[\frac{\sqrt{a}}{1+\sqrt{a}}*\frac{1-\sqrt{a}}{1-\sqrt{a}}=\frac{\sqrt{a}(1-\sqrt{a})}{1-a}=\frac{\sqrt{a}-a}{1-a}\]

hero (hero):

myininaya, I got the same answer you got....You just left before I could simplify it further

hero (hero):

My answer wasn't incorrect, just not simplified fully

myininaya (myininaya):

i didn't say it was. i just wanted to type it myself lol